A gas mixture with a total pressure of 755  mbar contains each of the following gases at the indicated partial pressures: CO2, 122  mbar ; Ar, 231  mbar ; and O2, 193  mbar . The mixture also contains helium gas.

What is the partial pressure of the helium gas? (in mBar)

What mass of helium gas is present in a 12.8  L sample of this mixture at 270  K ? (in g)
Answer #1

1)

p(He) = Total pressure - p(CO2) - p(Ar) - p(O2)

= 755 - 122 - 231 - 193

= 209 mbar

Answer: 209 mbar

2)

Given:

P = 209 mbar

= 0.209 bar

= (0.209/1.01325) atm

= 0.2063 atm

V = 12.8 L

T = 270.0 K

find number of moles using:

P * V = n*R*T

0.2063 atm * 12.8 L = n * 0.08206 atm.L/mol.K * 270 K

n = 0.1192 mol

Molar mass of He = 4.003 g/mol

use:

mass of He,

m = number of mol * molar mass

= 0.1192 mol * 4.003 g/mol

= 0.477 g

Answer: 0.477 g